OpenStudy (anonymous):

I'll give a Medal for the Correct Answer! What is the slope of the line that passes through the pair of points? (-5/3 - 1) & (-2, 9/2)

OpenStudy (anonymous):

WAit, what is the first coord?

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

coord?

OpenStudy (anonymous):

they meant coordinate

OpenStudy (anonymous):

slope = (y2-y1) / (x2-x1) = (9/2 +1)/ (-2+5) = (11/2) / (3) = 11/6

OpenStudy (anonymous):

wait i messed up

OpenStudy (anonymous):

(11/2) / ( -2 + (5/3) ) = (11/2) / ( -1/3) = ( -33/2)

OpenStudy (anonymous):

how did you get the 11/2?

OpenStudy (anonymous):

Here's what you do: 9/2--1 over -2--5 This is 11/2 over 3... so dived 11/2 by 3.. It equals 11/6, that's your slope

OpenStudy (anonymous):

You there? any questions?

OpenStudy (anonymous):

ohh ok yea that makes sense. But my slope can't be 11/6 because its not an option on the question.

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

2/33 -2/33 -33/2 33/2

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

oh i think I read it wrong before... lemme see what I get now.. sorry for the wait

OpenStudy (anonymous):

it alright

OpenStudy (anonymous):

I got -11/6..... maybe the answers are messed up? I don't know....

OpenStudy (anonymous):

Sorry... I know I did the slope formula right.... I don't know why it's not an option

OpenStudy (anonymous):

its alright, I appreciate for the effort though!